# Quantum thermalization and the route to ergodicity

### v1

@@ Abstract
Random matrix theory (RMT), and specifically the theory of Wigner banded random matrices (WBRM) [1, 2], have been proposed as a framework to describe ergodization and thermalization. Consider the quantum thermalization between coupled subsystems of a bipartite disordered system, where each subsystem by itself is generically chaotic. The Hamiltonian is ${H=H_{0}+\lambda V}$ where the inter-system interaction $\lambda V$ mixes the eigenstates of the uncoupled subsystems. It has been previously realized that generic perturbations of a quantized chaotic system (QCS) can be modeled by WBRM [3, 4, 5, 6, 7, 8, 9]. It follows that for relatively weak coupling the mixing of the levels is described by the Wigner Lorentzian, while for strong coupling the line-shape becomes non-perturbative. Namely, for WBRM it is a semicircle line-shape, while for a QCS it is a semiclassical line-shape that reflects the underlying phase-space geometry. The latter is implied by the association of chaotic eigenstates with micro-canonical distributions in Berry’s conjecture and the eigenstate thermalization hypotheis (ETH) [10, 11, 12, 13, 14, 15, 16].
As shown below, the application of WBRM theory to the thermalization of two chaotic subsystems is rather subtle. For illustration we consider a bi-partite Bose-Hubbard model (BHM) [17, 18, 19, 20, 21, 22, 23, 24]. The eigenstates of the disconnected subsystems $H_{0}|x,n\rangle=\varepsilon_{x,n}|x,n\rangle$ are direct products of subsystem eigenstates with well-defined integer subsystem occupations, indexed by two good quantum numbers: the occupation difference $x$ and the index $n$ of the state within the factorized spectrum (for a given $x$). These are plotted in Fig. 1, where each point represents an unperturbed eigenstate. The coupling $\left\langle x,n\middle|V\middle|x^{\prime},n^{\prime}\right\rangle$ connects only states with ${|x-x^{\prime}|=1}$. This 2D geometry of the basis for the BHM is very different from the 1D geometry of the WBRM (left panel of Fig. 1), with substantial implications on the route to ergodicity.
@@
Outline.– We introduce a critical discussion of the energy-shell concept. Then we explore the route towards ergodicity for the BHM and contrast it with the WBRM.
The BHM.– Consider the bipartite BHM for $N$ particles in $M=M_{A}{+}M_{B}$ sites, where $M_{\alpha}$ is the number of sites in subsystem $\alpha=A,B$. The Hamiltonian is,
@@
In the above, the boson operator ${\hat{a}}_{\alpha,i}$ annihilates a particle in the $i$-th site of the $\alpha$ subsystem with population ${\hat{n}}_{\alpha,i}={\hat{a}}_{\alpha,i}^{\dagger}{\hat{a}}_{\alpha,i}$, the intra-system hopping frequency is $K$, and $\langle i,j\rangle$ denotes nearest-neighbor pairs. The coupling strength is $U$, and the weak on-site random potential $\varepsilon_{\alpha,i}$ breaks parity. In the numerical demonstrations we have ${M_{A}{=}M_{B}{=}4}$, $N{=}10$, and ${|\varepsilon_{\alpha,i}|<0.05}$, while ${K{=}1}$, and $U{=}1$.
@@
where $D_{x}=D_{M_{A}}\left(\frac{N{+}x}{2}\right)\times D_{M_{B}}\left(\frac{N{-}x}{2}\right)$ is the dimension of the $x$ sector of the spectrum (i.e. the number of states in each column of the diagram in Fig. 1) and $x=-N,-N{+}2,\cdots,N{-}2,N$. In the absence of inter-system coupling ($\lambda=0$), the eigenenergies of the unperturbed Hamiltonian $H_{0}=H_{A}+H_{B}$ are $\varepsilon_{x,n}$, where the running index ${n=1,\dots,{{\color[rgb]{0.6,0,0.1}D_{x}}}}$ labels the $x$-sorted unperturbed energies.
Eigenstates of the coupled subsystems.– Once the coupling $\lambda$ it turned on, the subsystems can exchange particles, and the coordinate $x$ is no longer a good quantum number. Accordingly we use a running index ${\nu=1,\cdots,D_{M}(N)}$ to label the many-body eigenstates and the respective eigenenergies $E_{\nu}$. In the presentated results below, these energies are scaled as ${E:=(E{-}E_{\rm min})/(E_{\rm max}{-}E_{\rm min})}$, where $E_{\rm min}$ and $E_{\rm max}$ are the ground and top energies.
@@
$p_{x}^{(\nu)}$ $=$ $\ \sum_{n}p_{x,n}^{(\nu)}$ (5)
$p^{(\nu)}(\varepsilon)$ $=$ $\ \sum_{x,n}p_{x,n}^{(\nu)}\,\delta(\varepsilon-\varepsilon_{x,n})$ (6)
@@
$\mathcal{M}_{\nu}$ $=$ $\left[\sum_{x,n}\big{|}p_{x,n}^{(\nu)}\big{|}^{2}\right]^{-1}\!\!=\text{Participation number}$ (8)
@@
Figure 2: The width of the energy shell. (a) The width $\Delta x_{\text{shl}}$ versus the width $\delta E$ of the arbitrary energy window about ${E=0.41666}$ for $\lambda=10^{-3}$ (solid blue line), $10^{-2}$ (dashed red), $10^{-1}$ (dotted orange), and 1.01.01.0 (dash-dotted purple). The width of the energy shell is well defined (assuming quantum mechanically large but classically small $\delta E$) and the dependence of the obtained shell width on the coupling strength $\lambda$ is weak. (b,c) The dependence of $\Delta x_{\rm shl}$, ${\cal M}_{\rm shl}$ and $\Delta\varepsilon_{\rm shl}$ on the coupling $\lambda$. In the latter case one observe a residual value for small $\lambda$, corresponding to the arbitrary $\delta E$ (dashed line is for $\delta E=0.0333$ whereas dash-dotted line is for $\delta E=0.0166$). This numerical artifact can be eliminated by looking on ${\overline{\Delta\varepsilon_{\nu}}}$ (solid line), even if the eigenstates are non-ergodic! The dotted line is a guide to the eye, corresponding to linear $\Delta\varepsilon_{\rm shl}(\lambda)$ dependence
The energy shell.– Semiclassically, a microcanonical distribution occupies an energy shell of thickness $\delta E$ in phase-space. In the classical context the ‘energy surface’ limit $\delta E\rightarrow 0$, is well defined. Quantum mechanically, the thickness $\delta E$ should be large with respect to the mean level-spacing but small with respect to the energy scale on which this spacing varies, i.e. the shell width should be ‘classically small but quantum mechanically large’.
@@
By contrast, the energy dispersion $\Delta\varepsilon_{\rm shl}$ approaches the arbitrary window width $\delta E$ at small $\lambda$. This numerical artifact indicating that $\delta E$ is still too large. The dispersion $\Delta\varepsilon_{\rm shl}$ is the convolution of the individual widths $\Delta\varepsilon_{\nu}$ with the arbitrary $\delta E$. Physically meaningful result is obtained provided the contribution of $\delta E$ is negligible. It turns out that this numerical limitation can be avoided, because for classically small $\lambda$ we get [27]
@@
This holds even if the eigenstates are non-ergodic, as further discussed below under ‘False ergodicity measures’. Fig. 2c demonstrates the validity of this statement, and the implied linear dependence of $\Delta\varepsilon_{\rm shl}$ on $\lambda$.
Distinguishing ergodicity from chaos.– It is a common misconception to identify “quantum chaos” with ”quantum ergodicity”. Quantum chaos is identified by the spectral statistics. In Fig. 3 we display a re-scaled version of a common measure for level repulsion, namely the mean ratio of consecutive level spacings [25],
@@
The proper measure for the quantum ergodicity of the eigenstates, is the mean participation number, $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}$ (dashed line in Fig. 3). The $\mathcal{M}_{\nu}$ are calculated for the individual eigenstates, while $\mathcal{M}_{\rm shl}$ is calculated from the $p_{x,n}$ of the energy shell. For GOE statistics, a value of $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}\approx 1/3$ indicates the ergodicity of the eigenstates. The 1/31/31 / 3 factor emerges because the $p_{x,n}^{(\nu)}$ probabilities are not uniformly spread over the energy shell but statistically fluctuate.
@@
Comparing the two measures, we note that the ergodic regime in Fig. 3 is much smaller than the chaotic regime indicated by spectral statistics. For a substantial $\lambda$ range the system is chaotic, yet its eigenstates are not ergodic.
@@
The ETH ergodicity measures.– The ETH predicts that for ergodic eigenstates $|E_{\nu}\rangle$ the expectation value of any local observable, such as $\left\langle x\right\rangle_{\nu}$, will exhibit low state-to-state fluctuations, with dispersion ${\sigma_{x}\propto 1/\sqrt{\mathcal{M}}}$. The mere decrease of state-to-state fluctuations however, is not a sufficient indication for ergodicity, as clearly demonstrated by comparing the earlier (weak coupling) drop of $\sigma_{x}$ with the later (stronger coupling) rise of $\overline{{\cal M}_{\nu}}$ in Fig. 3.
@@
Figure 3: Chaos and Ergodicity Measures. Misc measures are calculated for the energy shell ${0.4<E<0.4333}$ as a function of the coupling strength $\lambda$, and normalized per the expected GOE values. The spectral measure for chaos $\tilde{r}$ (square markers) is correlated with $\overline{\Delta x_{\nu}}/\Delta x_{\rm shl}$ (solid line). The ETH measure $\sigma_{x}/\Delta x_{\rm shl}$ (dotted line) is an insufficient measure for ergodization. The measure $\overline{\mathcal{M}_{\nu}}/(\mathcal{M}_{\rm shl}/3)$ (dashed line) is an effective measure for the actual identification of the ergodic regime. We also plot the mean peak probability $\overline{P_{\nu}}$ (dash-dotted line), which is used to identify the first-order perturbative regime (arbitrarily defined as $\overline{P_{\nu}}>0.6$). Inverted triangles mark the coupling strengths for the states presented Fig. 4.
[…]

### v2

@@ Abstract
I Introduction
Random matrix theory (RMT), and specifically the theory of Wigner banded random matrices (WBRM) [1, 2], have been proposed as a framework to describe ergodization and thermalization. Consider the quantum thermalization between coupled subsystems of a bipartite disordered system, where each subsystem by itself is generically chaotic. The Hamiltonian is $H=H_{0}+\lambda V$} where the inter-system interaction $\lambda V$ mixes the eigenstates of the uncoupled subsystems.
It has been previously realized that generic perturbations of a quantized chaotic system (QCS) can be modeled by WBRM [3, 4, 5, 6, 7, 8, 9]. It follows that for relatively weak coupling the mixing of the levels is described by the Wigner Lorentzian, while for strong coupling the line-shape becomes non-perturbative. Namely, for WBRM it is a semicircle line-shape, while for a QCS it is a semiclassical line-shape that reflects the underlying phase-space geometry. The latter is implied by the association of chaotic eigenstates with micro-canonical distributions in Berry’s conjecture and the eigenstate thermalization hypotheis (ETH) [10, 11, 12, 13, 14, 15, 16].
Later studies have considered the route to ergodicity and thermalization of large arrays, with special emphasis on 1D chains [17, 18, 19, 20], utilizing both spectral and novel ETH measures [21, 22]. In such studies $H_{0}$ is integrable while $\lambda V$} spoils the integrability. Here, we consider a more appropriate modeling of thermalization, closer in spirit to the traditional thermodynamic concept, where Boltzmann’s ‘molecular chaos’ applies equally well to the constituent thermalizing subsystems. Accordingly, we consider below a model for thermalization, where $\lambda V$} is the coupling between the quantized chaotic subsystems. Each subsystem is chaotic by itself, while the coupling allows exchange of particles and energy. This configuration may be thought of as ‘vintage thermalization’ because it reflects the traditional 19th century reasoning of thermodynamics. Similar agenda has been presented in [23].
As shown below, the application of WBRM theory to vintage thermalization is rather subtle. For illustration we consider a bi-partite Bose-Hubbard model (BHM) [24, 25, 26, 27, 28, 29, 30, 31]. The eigenstates of the disconnected subsystems $H_{0}|x,n\rangle=\varepsilon_{x,n}|x,n\rangle$ are direct products of subsystem eigenstates with well-defined integer subsystem occupations, indexed by two good quantum numbers: the occupation difference $x$ and the index $n$ of the state within the factorized spectrum (for a given $x$). These are plotted in Fig. 1, where each point represents an unperturbed eigenstate. The coupling $\left\langle x,n\middle|V\middle|x^{\prime},n^{\prime}\right\rangle$ connects only states with $|x-x^{\prime}|=1$}. This 2D geometry of the basis for the BHM is very different from the 1D geometry of the WBRM (left panel of Fig. 1), with substantial implications on the route to ergodicity.
@@ I Introduction
Outline.– We introduce the model and basic definitions in Sections II and III. Critical discussion of the energy-shell concept is provided in Section IV. Identification of chaos and ergodicity are explained in Sections V and VI. Then we explore the route towards ergodicity for the BHM in Section VII, and contrast it with the WBRM. Auxiliary issues related to ergodicity measures and regime boundaries, are discussed in Sections VIII and Section IX. Finally, discussion and summary are provided in Section X. Appendix A provides a terse summary of the WBRM regimes, while Appendix B presents extra numerical demonstration for the irrelevance of subsystem geometry.
II The BHM
Consider the bipartite BHM for $N$ particles in $M=M_{A}{+}M_{B}$ sites, where $M_{\alpha}$ is the number of sites in subsystem $\alpha=A,B$. The Hamiltonian is,
@@ II The BHM
In the above, the boson operator ${\hat{a}}_{\alpha,i}$ annihilates a particle in the $i$-th site of the $\alpha$ subsystem with population ${\hat{n}}_{\alpha,i}={\hat{a}}_{\alpha,i}^{\dagger}{\hat{a}}_{\alpha,i}$, the intra-system hopping frequency is $K$, and $\langle i,j\rangle$ denotes nearest-neighbor pairs. The coupling strength is $U$, and the weak on-site random potential $\varepsilon_{\alpha,i}$ breaks parity. In the numerical demonstrations we have $M_{A}{=}M_{B}{=}4$}, N=10N{=}10, and $|\varepsilon_{\alpha,i}|<0.05$}, while K=1{K{=}1}, and U=1U{=}1.
@@
where $D_{x}=D_{M_{A}}\left(\frac{N{+}x}{2}\right)\times D_{M_{B}}\left(\frac{N{-}x}{2}\right)$ is the dimension of the $x$ sector of the spectrum (i.e. the number of states in each column of the diagram in Fig. 1) and $x=-N,-N{+}2,\cdots,N{-}2,N$. In the absence of inter-system coupling ($\lambda=0$), the eigenenergies of the unperturbed Hamiltonian $H_{0}=H_{A}+H_{B}$ are $\varepsilon_{x,n}$, where the running index $n=1,\dots,D_{x}$} labels the $x$-sorted unperturbed energies.
III Eigenstates of the coupled subsystems
Once the coupling $\lambda$ it turned on, the subsystems can exchange particles, and the coordinate $x$ is no longer a good quantum number. Accordingly we use a running index $\nu=1,\cdots,D_{M}(N)$} to label the many-body eigenstates and the respective eigenenergies $E_{\nu}$. In the presentated results below, these energies are scaled as $E:=(E{-}E_{\rm min})/(E_{\rm max}{-}E_{\rm min})$}, where $E_{\rm min}$ and $E_{\rm max}$ are the ground and top energies.
@@ III Eigenstates of the coupled subsystems
$p_{x}^{(\nu)}\$$=$ $\ \sum_{n}p_{x,n}^{(\nu)}$ (5)
$p^{(\nu)}(\varepsilon)\$$=$ $\ \sum_{x,n}p_{x,n}^{(\nu)}\,\delta(\varepsilon-\varepsilon_{x,n})$ (6)
@@
$\mathcal{M}_{\nu}$ $=$ $\left[\sum_{x,n}\big|p_{x,n}^{(\nu)}\big|^{2}\right]^{-1}\!\!=\text{Participation number}\ \ \ \ \$(8)
@@
Figure 2: The width of the energy shell. (a) The width $\Delta x_{\text{shl}}$ versus the width $\delta E$ of the arbitrary energy window about E=0.41666{E=0.41666} for $\lambda=10^{-3}$ (solid blue line), $10^{-2}$ (dashed red), $10^{-1}$ (dotted orange), and $1.0$ (dash-dotted purple). The width of the energy shell is well defined (assuming quantum mechanically large but classically small $\delta E$) and the dependence of the obtained shell width on the coupling strength $\lambda$ is weak. (b,c) The dependence of $\Delta x_{\rm shl}$, ${\cal M}_{\rm shl}$ and $\Delta\varepsilon_{\rm shl}$ on the coupling $\lambda$. In the latter case one observe a residual value for small $\lambda$, corresponding to the arbitrary $\delta E$ (dashed line is for $\delta E=0.0333$ whereas dash-dotted line is for $\delta E=0.0166$). This numerical artifact can be eliminated by looking on $\overline{\Delta\varepsilon_{\nu}}$} (solid line), even if the eigenstates are non-ergodic! The dotted line is a guide to the eye, corresponding to linear $\Delta\varepsilon_{\rm shl}(\lambda)$ dependence
IV The energy shell
Semiclassically, a microcanonical distribution occupies an energy shell of thickness $\delta E$ in phase-space. In the classical context the ‘energy surface’ limit $\delta E\rightarrow 0$, is well defined. Quantum mechanically, the thickness $\delta E$ should be large with respect to the mean level-spacing but small with respect to the energy scale on which this spacing varies, i.e. the shell width should be ‘classically small but quantum mechanically large’.
@@ IV The energy shell
By contrast, the energy dispersion $\Delta\varepsilon_{\rm shl}$ approaches the arbitrary window width $\delta E$ at small $\lambda$. This numerical artifact indicating that $\delta E$ is still too large. The dispersion $\Delta\varepsilon_{\rm shl}$ is the convolution of the individual widths $\Delta\varepsilon_{\nu}$ with the arbitrary $\delta E$. Physically meaningful result is obtained provided the contribution of $\delta E$ is negligible. It turns out that this numerical limitation can be avoided, because for classically small $\lambda$ we get
@@
This holds even if the eigenstates are non-ergodic, as further discussed in the following subsections with consequences that are highlighted in Sec. (IX). Fig. 2c demonstrates the validity of this statement, and the implied linear dependence of $\Delta\varepsilon_{\rm shl}$ on $\lambda$.
IV.1 Width of the energy shell
Denoting unperturbed and exact eigenstates as $|\varepsilon\rangle$ and $|E\rangle$ respectively, the local density of states (LDOS) and the inverse-LDOS (the lineshape of the eigenstats) are defined respectively as follows:
$P^{(\varepsilon)}(E)$ $=$ $|\left\langle E\middle|\varepsilon\right\rangle|^{2}$ (12)
$p^{(E)}(\varepsilon)$ $=$ $|\left\langle\varepsilon\middle|E\right\rangle|^{2}$ (13)
What we call $\Delta\varepsilon_{\rm shl}$ is the $\varepsilon$-width of $p^{(E)}(\varepsilon)$, while $\Delta_{E}$ is defined as the $E$-width of $P^{(\varepsilon)}(E)$}. Let us first clarify these definitions in a purely classical context. Assume that $H(X,P)=H_{0}+\lambda V$}. Consider a microcanonical distribution $H_{0}(X,P)=\varepsilon$} and define E=H(X,P){E=H(X,P)}. We have the trivial identity
$\Delta_{E}^{2}=\lambda^{2}\left[\left\langle V(X,P)^{2}\right\rangle_{\varepsilon}-\left\langle V(X,P)\right\rangle_{\varepsilon}^{2}\right]$ (14)
where the the $\varepsilon$ subscript indicates an $H_{0}$-microcanconical phase-space average. This result implies strict linear dependence $\Delta_{E}\propto\lambda$}. The dual statement for the inverse-LDOS is as follows. Consider a microcanonical distribution H(X,P)=E{H(X,P)=E} and define $\varepsilon=H_{0}(X,P)$}. We have the trivial identity
$\Delta\varepsilon^{2}=\lambda^{2}\left[\left\langle V(X,P)^{2}\right\rangle_{E}-\left\langle V(X,P)\right\rangle_{E}^{2}\right]$ (15)
where the the $E$ subscript indicates an $H$-microcanconical phase-space average. Here there is an implicit $\lambda$-dependence of the latter. But if the difference between the surfaces H(X,P)=E{H(X,P)=E} and the surfaces $H_{0}(X,P)=\varepsilon$} is small, the $H$ microcanonical average can be calculated with an unperturbed surface of $H_{0}$, and the same result will be obtained, approximately, provided $\lambda$ is classically small enough.
IV.2 Second moment energy spreading
Here we provide the proof that the second-moment of the energy spreading has a robust linear dependence of $\lambda$, provided the perturbation is classically small, irrespective of quantum ergodicity. The proof is merely a translation of the classical statement Eq. (15) into quantum language. We assume that $H=H_{0}+\lambda V$}. Let $\left|\varepsilon\right\rangle$ be the eigenvalues of $H_{0}$, and let $\left|E\right\rangle$ be a chosen eigenstate of $H$. Then it follows that
$\Delta\varepsilon\ \$$=$ $\ \ \lambda\sqrt{\left\langle E\middle|V^{2}\middle|E\right\rangle-\left\langle E\middle|V\middle|E\right\rangle^{2}}$ (16)
$\approx$ $\ \ \lambda\sqrt{[V^{2}]_{\varepsilon,\varepsilon}-[V]_{\varepsilon,\varepsilon}^{2}}\ \ \equiv\ \ \Delta_{E}$ (17)
with $\varepsilon=E$}. This result, which we further explain below, is independent of whether the eigenstate is quantum ergodic or not. The validity of this observation can be optionally verified by direct calculation of the second moment spreading for the non-ergodic perturbative eigenstates of Eq. (20).
Let us elaborate on the two steps in Eq. (16). The first equality is based on a trivial identity that holds for the k=1,2{k=1,2} moments (and cannot be extended to higher moments due to operator ordering issue):
$\left\langle\varepsilon^{k}\right\rangle$ $=$ $\sum_{\varepsilon}|\left\langle\varepsilon\middle|E\right\rangle|^{2}\varepsilon^{k}\ \ =\ \ \left\langle E\middle|H_{0}^{k}\middle|E\right\rangle$
$=$ $\left\langle E\middle|(H-\lambda V)^{k}\middle|E\right\rangle$
$=$ $\left\{\begin{matrix}E-\lambda\left\langle E\middle|V\middle|E\right\rangle&|\text{for}\ k{=}1\\ E^{2}-2\lambda E\left\langle E\middle|V\middle|E\right\rangle+\left\langle E\middle|V^{2}\middle|E\right\rangle&|\text{for}\ k{=}2\end{matrix}\ \ \ \ \ \right$.
The second equality in Eq. (16) holds for a classically small $\lambda$, because the matrix elements of any generic operator between perturbed eigenstates is approximately equal to the matrix elements between the unperturbed eigenstates, namely
$\left\langle E\middle|A\middle|E\right\rangle\ \approx\ \left\langle\varepsilon\middle|A\middle|\varepsilon\right\rangle\ \equiv\ A_{\varepsilon,\varepsilon}\ \ \text{with}\ \varepsilon\approx E$ (18)
This approximation is based on the observation that $\left\langle\psi\middle|A\middle|\psi\right\rangle=\mbox{trace}[A\rho]$} with $\rho=\left|\Psi\right\rangle\left\langle\Psi\right|$} can be written as a phase-space integral over Wigner functions. If the microcanonical energy surface that supports $\rho$ is not related to the phase-space contours of $A$, the trace overlap is rather robust. A small perturbation to $\psi$ will not affect systematically the result (i.e. disregarding quantum fluctuations).
V Distinguishing ergodicity from chaos
It is a common misconception to identify “quantum chaos” with ”quantum ergodicity”. Quantum chaos is identified by the spectral statistics. In Fig. 3 we display a re-scaled version of a common measure for level repulsion, namely the mean ratio of consecutive level spacings [32],
@@ V Distinguishing ergodicity from chaos
The proper measure for the quantum ergodicity of the eigenstates, is the mean participation number, $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}$ (dashed line in Fig. 3). The $\mathcal{M}_{\nu}$ are calculated for the individual eigenstates, while $\mathcal{M}_{\rm shl}$ is calculated from the $p_{x,n}$ of the energy shell. For GOE statistics, a value of $\overline{\mathcal{M}_{\nu}}/\mathcal{M}_{\rm shl}\approx 1/3$ indicates the ergodicity of the eigenstates. The $1/3$ factor emerges because the $p_{x,n}^{(\nu)}$ probabilities are not uniformly spread over the energy shell but statistically fluctuate.
@@
Comparing the two measures, we note that the ergodic regime in Fig. 3 is much smaller than the chaotic regime indicated by spectral statistics. For a substantial $\lambda$ range the system is chaotic, yet its eigenstates are not ergodic. Similar discrepancies between chaos and ergodization measures have been previously observed without further exploration in [33, 34, 35, 36]
VI The ETH ergodicity measures
@@
The ETH predicts that for ergodic eigenstates $|E_{\nu}\rangle$ the expectation value of any local observable, such as $\left\langle x\right\rangle_{\nu}$, will exhibit low state-to-state fluctuations, with dispersion $\sigma_{x}\propto 1/\sqrt{\mathcal{M}}$}. The mere decrease of state-to-state fluctuations however, is not a sufficient indication for ergodicity, as clearly demonstrated by comparing the earlier (weak coupling) drop of $\sigma_{x}$ with the later (stronger coupling) rise of $\overline{{\cal M}_{\nu}}$ in Fig. 3.
@@
Figure 3: Chaos and Ergodicity Measures. Misc measures are calculated for the energy shell 0.4<E<0.4333{0.4<E<0.4333} as a function of the coupling strength $\lambda$, and normalized per the expected GOE values. The spectral measure for chaos $\tilde{r}$ (square markers) is correlated with $\overline{\Delta x_{\nu}}/\Delta x_{\rm shl}$ (solid line). The ETH measure $\sigma_{x}/\Delta x_{\rm shl}$ (dotted line) is an insufficient measure for ergodization. The measure $\overline{\mathcal{M}_{\nu}}/(\mathcal{M}_{\rm shl}/3)$ (dashed line) is an effective measure for the actual identification of the ergodic regime. We also plot the mean peak probability $\overline{P_{\nu}}$ (dash-dotted line), which is used to identify the first-order perturbative regime (arbitrarily defined as $\overline{P_{\nu}}>0.6$). Inverted triangles mark the coupling strengths for the states presented Fig. 4.
Figure 4: Representative eigenstates. Upper row: Images of $p^{(\nu)}_{x,n}$ for a representative eigenstate with $E\approx 0.425$}. Each panel is for a different coupling corresponding (from left to right) to the Perturbative, Mixing, Wigner, and Ergodic regimes. Second row: The resulting imbalance distribution $p_{x}^{(\nu)}$. Third row: The $p_{x,n}^{(\nu)}$ probabilities are plotted versus $(x,n)$, arranged in order of ascending unperturbed energy. Bottom row: The same $p_{x,n}^{(\nu)}$ probabilities are compared on a log scale to the coarse-grained 1st-order perturbation profiles (solid lines). White lines are coarse-graining of $p_{x,n}^{(\nu)}$.
VII Contrast with Wigner theory